OpenStudy (anonymous):

What is the solution set for 6(3+4x)-6 >= 204

OpenStudy (anonymous):

the solution is found by following these steps: 6(3+4x)-6 >= 204 add 6 to both sides 6(3+4x) >= 210 divide both sides by 6 3+4x >= 210/6 subtract 3 from both sides 4x >= 210/6 - 3 find a common denominator for 210/6 and 3 and subtract 210/6 - 18/6 = 192 4x >= 192 divide both sides by 4 x >= 48

OpenStudy (anonymous):

err that should be 192/6 so 4x >= 192/6 4x>= 32 x>=8 sorry about that ;)

OpenStudy (anonymous):

so it would be >= 8